the perimeter of the rectangular garden is 60 feet. if the length is twice the width what are the dimensions of the rectangular garden
Answer by KlambinRonsentaswer(36)

I'm glad to help you. :)
let the wigth =x
then the length =2x

You can get
2·(2x+x)=60
2·(2x+x)÷2=60÷2
2x+x=30
3x=30
x=10

So 2x=2×10=20

The width of the garden is 10 feet, the length is 20 feet and the area is 200 square feet.